### v1

# The odd independence number of graphs, I: Foundations and classical classes

Abstract
An odd independent set $S$ in a graph $G=(V,E)$ is an independent set of vertices such that, for every vertex $v\in V\,\diagdown\,S$, either $N(v)\cap S=\varnothing$ or $|N(v)\cap S|\equiv 1$ (mod 2), where $N(v)$ stands for the open neighborhood of $v$. The largest cardinality of odd independent sets of a graph $G$, denoted $\alpha_{\mathrm{od}}(G)$, is called the odd independence number of $G$.
This new parameter is a natural companion to the recently introduced strong odd chromatic number. A proper vertex coloring of a graph $G$ is a strong odd coloring if, for every vertex $v\in V(G)$, each color used in the neighborhood of $v$ appears an odd number of times in $N(v)$. The minimum number of colors in a strong odd coloring of $G$ is denoted by $\chi_{\mathrm{so}}(G)$.
A simple relation involving these two parameters and the order $|G|$ of $G$ is $\alpha_{\mathrm{od}}(G)\cdot\chi_{\mathrm{so}}(G)\geq|G|$, parallel to the same on chromatic number and independence number.
We develop several basic inequalities concerning $\alpha_{\mathrm{od}}(G)$, and use already existing results on strong odd coloring, to derive lower bounds for odd independence in many families of graphs.
We prove that $\alpha_{\mathrm{od}}(G)=\alpha(G^{2})$ holds for all claw-free graphs $G$, and present many results, using various techniques, concerning the odd independence number of cycles, paths, Moore graphs, Kneser graphs, the complete subdivision $S(K_{n})$ of $K_{n}$, the half graphs $H_{n,n}$, and $K_{p}\,\Box\,K_{q}$. Further, we consider the odd independence number of the hypercube $Q_{d}$ and also of the complements of triangle-free graphs. Many open problems for future research are stated.
Keywords: independence number, odd independence number, strong odd coloring.
AMS Subject Classification 2020: 05C15, 05C69 (primary), 05C76 (secondary)
1 Introduction
An odd independent set $S$ in a graph $G=(V,E)$ is an independent set of vertices such that, for every vertex $v\in V\,\diagdown\,S$, either $N(v)\cap S=\varnothing$ or $|N(v)\cap S|\equiv 1$ (mod 2), where $N(v)$ stands for the open neighbourhood of $v$. The largest cardinality of odd independent sets in $G$, denoted $\alpha_{\mathrm{od}}(G)$, is called the odd independence number of $G$. This new parameter is a natural companion to the recently introduced strong odd chromatic number [3, 7, 9, 10, 11]. A proper vertex coloring of a graph $G$ is a strong odd coloring if, for every vertex $v\in V(G)$, each color used in the neighborhood of $v$ appears an odd number of times in $N(v)$. The minimum number of colors in a strong odd coloring of $G$ is denoted by $\chi_{\mathrm{so}}(G)$. This paper is aimed at introducing and starting a systematic study of the odd independence number.
In Section 2 we present basic inequalities relating $\alpha_{\mathrm{od}}(G)$, $\chi_{\mathrm{so}}(G)$ and other graphs parameters like $|G|$ (the order of $G$) and $\Delta(G)$ (the maximum degree of $G$), similar to such classical inequalities involving $\alpha(G)$ and $\chi(G)$. We also supply upper bounds for $\alpha_{\mathrm{od}}(G)$ for $d$-regular graphs when $d$ is even, or when $G$ is a $K_{1,r}$-free graph, or when any two adjacent vertices in $G$ have at least $\lambda$ vertices in common. We also mention that $\alpha_{\mathrm{od}}(G)=\alpha(G)$ holds for every regular bipartite graph $G$ with odd degree, and we further prove $\alpha_{\mathrm{od}}(G)=\alpha(G^{2})$ for claw-free graphs $G$, complementing the result [3] that $\chi_{\mathrm{so}}(G)=\chi(G^{2})$ holds for claw-free graphs. Then we present various lower bounds concerning $\alpha_{\mathrm{od}}(G)$ using known results on $\chi_{\mathrm{so}}(G)$ and the inequality $\alpha_{\mathrm{od}}(G)\geq|G|/\chi_{\mathrm{so}}(G)$, which serves as a benchmark for further results.
In Section 3 we develop estimates concerning $\alpha_{\mathrm{od}}(G)$ for many special classes of graphs, including the Petersen graph, Hoffman–Singleton graph and Kneser graphs, helping to assimilate the notion of odd independence.
In Section 4 we present methods to recursively construct graphs from given ones, by means of graph operations inspired by the Cartesian product of graphs, which allow us to conclude a lower bound on the odd independence number of the produced graph in terms of the building blocks.
In Section 5 we apply the constructions introduced in Section 4 to compute a lower bound for $\alpha_{\mathrm{od}}(Q_{d})$, the $d$-dimensional cube, when $d$ is even. The case of $d$ odd is trivially covered by the simple fact that for every $d$-regular bipartite graph $G$ with $d$ odd, $\alpha_{\mathrm{od}}(G)=\alpha(G)=|G|/2$. In contrast, the case when $d$ is even is challenging.
In Section 6 we consider complements $\overline{G}$ of triangle-free graphs $G$. These complementary graphs have $\alpha(\overline{G})\leq 2$, hence also $\alpha_{\mathrm{od}}(\overline{G})\leq 2$. We analyse the possible values of $\alpha_{\mathrm{od}}(G)$ with respect to the diameter of $G$ and $\overline{G}$.
In Section 7 we improve upon one of the basic inequalities presented in Section 2, namely $\alpha_{\mathrm{od}}(G)\geq|G|/(\Delta^{2}(G)+1)$. Using our results concerning Moore graphs, we prove that the stronger inequality $\alpha_{\mathrm{od}}(G)\geq|G|/(\Delta^{2}(G)-1)$ is valid for all $\Delta\geq 3$.
In Section 8 we list many open problems for future research, covering structural problems as well as computational complexity problems.
1.1 Notation
Let $G$ be any graph (simple, undirected). We denote its number $|V(G)|$ of vertices by $|G|$ (the order of $G$), and its number of edges by $e(G)$. As usual, $\alpha(G)$ and $\chi(G)$ denote the independence number and the chromatic number of $G$, respectively. Further standard notations are $N(v)$ (open neighborhood), $N[v]$ (closed neighborhood), and $\deg_{G}(v)$ or $\deg(v)$ (degree) of a vertex $v\in V(G)$, and similarly $N(X)$ and $N[X]$ for the corresponding neighborhoods (open/closed) of a set $X\subset V(G)$. The average degree of graph $G$ is $d(G):=\frac{1}{|G|}\sum_{v\in V(G)}\deg(v)=\frac{2e(G)}{|G|}$. In the square $G^{2}$ of $G$, two vertices are adjacent if and only if they are at distance at most 2 apart in $G$ (i.e., they are adjacent or have a common neighbor).
Well-known basic types of graphs are the complete graph $K_{n}$, the empty (edgeless) graph $E_{n}$, and the path $P_{n}$ of order $n$. The graph operations $G\cup H$ (vertex-disjoint union), $G+H$ (complete join), $tG$ (the vertex-disjoint union of $t$ copies of $G$) are also standard.
Beside $\alpha_{\mathrm{od}}(G)$ and $\chi_{\mathrm{so}}(G)$ we shall occasionally mention the further parameter $\chi_{\mathrm{o}}(G)$, termed odd chromatic number, which is the smallest number $k$ of colors admitting a proper vertex $k$-coloring such that, for each non-isolated vertex $v\in V(G)$, some color occurs an odd number of times in $N(v)$.
2 Some basic facts
In the next proposition we collect the basic relations between $\alpha_{\mathrm{od}}(G)$ and $\chi_{\mathrm{so}}(G)$, parallel to the classical relation between $\alpha(G)$ and $\chi(G)$.
Proposition 1.
Let $G=(V,E)$ be a graph on $n$ vertices.
  • $(1)$
    $\chi_{\mathrm{o}}(G)\leq\chi_{\mathrm{so}}(G)\leq\chi(G^{2})\leq\Delta(G^{2})+1\leq(\Delta(G))^{2}+1$.
  • $(2)$
    $\chi_{\mathrm{so}}(G)=\chi(G^{2})$ whenever $G$ is claw-free.
  • $(3)$
    $\alpha_{\mathrm{od}}(G)\geq n/\chi_{\mathrm{so}}(G)\geq n/\chi(G^{2})\geq n/(\Delta(G^{2})+1)\geq n/((\Delta(G))^{2}+1)$.
  • $(4.1)$
    $(n+1)^{2}\!/4\geq\alpha_{\mathrm{od}}(G)\chi_{\mathrm{so}}(G)\geq n$, and both bounds are best possible.
  • $(4.2)$
    $n+1\geq\alpha_{\mathrm{od}}(G)+\chi_{\mathrm{so}}(G)\geq 2\sqrt{n}$, and both bounds are best possible.
  • $(5)$
    $\alpha_{\mathrm{od}}(G)\geq\alpha(G^{2})$.
  • $(6)$
    $\alpha_{\mathrm{od}}(G)\geq\sum_{v\in V}1/(\deg_{G^{2}}(v)+1)\geq n/(d(G^{2})+1)$
    $\geq n/(d(G)\Delta(G)+1)\geq n/((\Delta(G))^{2}+1)$.
Proof.
(1) The only inequality that needs a proof is $\chi_{\mathrm{so}}(G)\leq\chi(G^{2})$. Let us consider a coloring of $G^{2}$ with $\chi(G^{2})$ colors. Then in $G$ every vertex $v$ has all the vertices of $N(v)$ getting distinct colors (as $N[v]$ forms a clique in $G^{2}$), hence any color present in $N(v)$ appears once.
(2) This was proved already in the paper on strong odd coloring, [3, Proposition 3].
(3) If $S_{1}\cup\cdots\cup S_{k}=V(G)$ is a partition into $k:=\chi_{\mathrm{so}}(G)$ odd independent sets, then $|G|=|S_{1}|+\dots+|S_{k}|\leq k\cdot\alpha_{\mathrm{od}}(G)=\alpha_{\mathrm{od}}(G)\cdot\chi_{\mathrm{so}}(G)$. This implies the leftmost inequality. The rest of the chain of inequalities follows from item (1).
(4.1)–(4.2) Clearly from (3) we have $\alpha_{\mathrm{od}}(G)\chi_{\mathrm{so}}(G)\geq n$. By the inequality between arithmetic and geometric means, denoting $\alpha_{\mathrm{od}}(G)=x$ and $\chi_{\mathrm{so}}(G)=y$, we obtain $x+y\geq 2\sqrt{xy}\geq 2\sqrt{n}$, proving the right side of (4.2). Now choose a maximum odd independent set $S$, color all its vertices with the same color, and the remaining n|S|n-|S| vertices with distinct colors, altogether using $n-\alpha_{\mathrm{od}}(G)+1$ colors. This is easily seen to be a strong odd coloring. Hence $n+1\geq\alpha_{\mathrm{od}}(G)+\chi_{\mathrm{so}}(G)$. Lastly, again by the inequality between arithmetic and geometric means, with the same notation as above, $(n+1)/2\geq(x+y)/2\geq\sqrt{xy}$ holds, hence we conclude that $(n+1)^{2}\!/4\geq\alpha_{\mathrm{od}}(G)\chi_{\mathrm{so}}(G)$.
Tightness of the lower bounds is shown by the graph $G=tK_{r}$, $n=tr$. Then $\chi_{\mathrm{so}}(G)=r$ and $\alpha_{\mathrm{od}}(G)=t$, hence $n=\alpha_{\mathrm{od}}(G)\chi_{\mathrm{so}}(G)$. In particular, for $t=r$ and $n=r^{2}$ we obtain $\alpha_{\mathrm{od}}(G)+\chi_{\mathrm{so}}(G)=2r=2\sqrt{n}$.
Tightness of the upper bounds is shown by the graph $G=K_{r}\cup tK_{1}$, $n=r+t$. Then $\chi_{\mathrm{so}}(G)=r$ and $\alpha_{\mathrm{od}}(G)=t+1$, hence $n+1=\alpha_{\mathrm{od}}(G)+\chi_{\mathrm{so}}(G)$. In particular, if $n=2k+1$ is odd, we take $r=k+1=(n+1)/2$ and $t=k$, hence obtain $\alpha_{\mathrm{od}}(G)\chi_{\mathrm{so}}(G)=r\cdot(t+1)=\frac{(n+1)^{2}}{4}$. If $n=2k$ is even, we can take $r=t=k$ and obtain $\alpha_{\mathrm{od}}(G)\chi_{\mathrm{so}}(G)=k^{2}+k=(k+1/2)^{2}-1/4=\lfloor\frac{(n+1)^{2}}{4}\rfloor$.
(5) Let $S$ be a maximum independent set in $G^{2}$ and consider $S$ as an independent set in $G$ (clearly it is). Then a vertex in $V\,\diagdown\,S$ is adjacent to at most one vertex in $S$, since if it was adjacent to two vertices $x,y\in S$ then $x,y$ are not independent in $G^{2}$ and thus cannot be both in $S$.
(6) Observe that $\deg_{G^{2}}(v)\leq\deg(v)+\deg(v)(\Delta(G)-1)=\deg(v)\Delta(G)$ holds for every vertex $v$. Hence, applying (5) and the Caro–Wei inequality, $\alpha_{\mathrm{od}}(G)\geq\alpha(G^{2})\geq\sum_{v\in V}1/(\deg_{G^{2}}(v)+1)\geq\sum_{v\in V}1/(\deg_{G}(v)\Delta(G)+1)\geq n/(d(G)\Delta(G)+1)$, the last step obtained via Jensen’s inequality.
Alternatively,
$d(G^{2})=2e(G^{2})/n\leq\bigl(\sum_{v\in V}\deg_{G^{2}}(v)\bigr)/n\leq\Delta(G)\bigl(\sum_{v\in V}\deg(v)\bigr)/n=\Delta(G)d(G)$
and, by Turán’s theorem,
$\alpha_{\mathrm{od}}(G)\geq\alpha(G^{2})\geq n/(d(G^{2})+1)\geq n/(d(G)\Delta(G)+1)\geq n/(\Delta^{2}(G)+1)\,.$
$\Box$
Proposition 2.
If $G$ is a bipartite graph with all degrees odd, then
  • $(i)$
    $\chi_{\mathrm{so}}(G)=2$ and $\alpha_{\mathrm{od}}(G)\geq|G|/2\geq(\alpha(G)+1)/2$;
  • $(ii)$
    if $G$ is also $d$-regular, with $d$ odd, then $\alpha_{\mathrm{od}}(G)=\alpha(G)=|G|/2$.
Proof.
By the degree assumption, each vertex class is an odd independent set, and one of them contains at least half of the vertices, hence proving $(i)$. (The lower bound $(\alpha(G)+1)/2$ on $|G|/2$ is valid since $G$ cannot be edgeless. Tightness is shown by the stars $K_{1,n-1}$, $n$ even.)
If $G$ is regular and bipartite, then it contains a perfect matching, consequently $\alpha(G)$ equals half the order of the graph. This proves $(ii)$. $\Box$
For $d$-regular graphs with $d$ even, we have the following upper bound.
Proposition 3.
Assume that $d$ is even. If $G$ is a $d$-regular graph, then $\alpha_{\mathrm{od}}(G)\leq\frac{d-1}{2d-1}\cdot|G|$, and this upper bound is tight for every $d\geq 2$ in the class of $d$-regular connected graphs of any order $n$ of the form $n=t\cdot(2d-1)$. Moreover, the bound is tight also in the class of $d$-regular bipartite graphs.
Proof.
Upper bound: Let $A$ be an odd independent set of size $|A|=\alpha_{\mathrm{od}}(G)$, and let $B$ be the set of vertices having at least one neighbor in $A$. Consider the bipartite graph $H$ formed by the AABB edges. Every $u\in A$ has degree $d$ in $H$ as $A$ is independent, and every $v\in B$ has degree at most d1d-1 in $H$, because it must be odd and $d$ is even. Thus, $d\cdot|A|\leq(d-1)\cdot|B|$, and consequently $|G|\geq|A|+|B|\geq|A|+\frac{d}{d-1}\cdot|A|$, implying the assertion.
Lower bound: Consider $t\geq 2$ copies of $K_{d,d-1}$, and take a perfect matching in the union of the $t$ $d$-element vertex classes in such a way that the obtained graph is connected. (If $t=1$, just take a matching of $d/2$ edges inside the vertex class of size $d$.) In particular, for $d=2$ this means $C_{3t}$.
To obtain bipartite examples, let $t=2$ and take the mentioned perfect matching between the two $d$-element vertex classes, keeping those two classes as independent sets. $\Box$
We can also state an asymptotic version of the above assertion, for graphs that are not far from being regular.
Proposition 4.
Let $\Delta$ be an even natural number, and for $j=1,2,\dots$ let $G_{j}$ be a graph on $n_{j}$ vertices, $n_{j}\to\infty$, of maximum degree $\Delta$ in which the number $z_{j}$ of vertices of degrees smaller than $\Delta$ satisfies $z_{j}=o(n_{j})$. Then $\limsup\alpha_{\mathrm{od}}(G_{j})/n\leq\frac{\Delta-1}{2\Delta-1}\,n$.
Proof.
Let $B_{j}$ be a maximum odd independent set of $G_{j}$, $Z_{j}\subset V(G_{j})$ the set of vertices of degree smaller than $\Delta$, and $B=B_{j}\,\diagdown\,Z_{j}$. Clearly, $|B_{j}|\geq n/(\Delta^{2}+1)$ and if $n_{j}$ is large enough then $B\neq\varnothing$ as $z_{j}=o(n_{j})$. Due to the parity conditions, every vertex not in $B_{j}$ is adjacent to at most $\Delta-1$ vertices in $B$. The number of edges emanating from $B_{j}$, say $e_{j}$, can be estimated as
$\Delta\cdot|B|=\Delta\cdot(|B_{j}|-z_{j})\leq e_{j}\leq(\Delta-1)(n_{j}-|B_{j}|)=(\Delta-1)(n_{j}-|B|-z_{j})\,.$
This gives $(2\Delta-1)|B|\leq(\Delta-1)(n_{j}-z_{j})$ and $\alpha_{\mathrm{od}}(G_{j})\leq|B|+z_{j}\leq\frac{\Delta-1}{2\Delta-1}\,(n_{j}-o(n_{j}))+o(n_{j})$. Hence $\limsup\alpha_{\mathrm{od}}(G_{j})/n\leq\frac{\Delta-1}{2\Delta-1}\,n$. $\Box$
Another extension, inspired by strongly regular graphs, can be stated as follows.
Proposition 5.
Let $G$ be a $\Delta$-regular graph on $n$ vertices, in which every two adjacent vertices have at least $\lambda$ common neighbors.
  • $(i)$
    If $\Delta-\lambda\equiv 0\ (\mathrm{mod}\ 2)$, then $\alpha_{\mathrm{od}}(G)\leq\frac{\Delta-\lambda-1}{2\Delta-\lambda-1}\,n$.
  • $(ii)$
    If $\Delta-\lambda=1\ (\mathrm{mod}\ 2)$, then $\alpha_{\mathrm{od}}(G)\leq\frac{\Delta-\lambda}{2\Delta-\lambda}\,n$.
Proof.
Let $B$ be a maximum odd independent set in $G$.
$(i)$ Consider $v\in V\,\diagdown\,B$. Either $v$ is not adjacent to any vertex in $B$, or it has at least one neighbor $u\in B$. In the latter case none of the (at least) $\lambda$ common neighbors of $u$ and $v$ can be in $B$ (as they are neighbors of $u$). So, $v$ can be adjacent to at most $\Delta-\lambda$ vertices in $B$, but in fact this means only $\Delta-\lambda-1$ because $\Delta-\lambda$ is even. Hence $v$ is adjacent to at most $\Delta-\lambda-1$ vertices in $B$.
Now count the number of edges between $B$ and $V\,\diagdown\,B$ in two ways and find $\Delta\cdot|B|\leq(\Delta-\lambda-1)(n-|B|)$. Hence $(2\Delta-\lambda-1)|B|\leq(\Delta-\lambda-1)n$, therefore $\alpha_{\mathrm{od}}(G)=|B|\leq\frac{\Delta-\lambda-1}{2\Delta-\lambda-1}\,n$, as claimed.
$(ii)$ The same proof applies, observing that in this case $v$ can be adjacent to at most $\Delta-\lambda$ members of $B$. $\Box$
The particular case $\lambda=0$ of $(ii)$ yields the upper bound $n/2$, which is tight e.g. in odd regular bipartite graphs.
Similar upper bounds concerning $K_{1,r}$-free graphs can be formulated. We give them in detail in the companion part II of this paper [4] where it is used for various grids and chessboard graphs, which are $K_{1,5}$-free.
We also observe that graphs for all feasible combinations of the parameters exist.
Proposition 6.
Graphs $G$ of order $n$ with $\alpha_{\mathrm{od}}(G)=\alpha(G)=k$ exist under any of the conditions below:
  • $(i)$
    $1\leq k\leq n$;
  • $(ii)$
    $1\leq k\leq n$, $k$ odd, $G$ connected;
  • $(iii)$
    $k\mid n$ with $k$ odd, $G$ regular and connected;
  • $(iv)$
    $2\leq k\leq n-2$, $k$ even, $G$ connected.
Proof.
For the above cases, we can take the following graphs.
$(i)$ Take $G=K_{n-k+1}\cup(k-1)K_{1}$, or more generally any graph consisting of $k$ complete components.
$(ii)$ The graph $G=K_{n-k}+E_{k}$, the split graph with complete join works; and for $k<n$ also any complete $t$-partite graph $K_{a_{1},\dots,a_{t}}$ where $a_{1}\geq a_{2}\geq\dots\geq a_{t}$, with $\sum a_{j}=n$ and $a_{1}=k$ odd.
$(iii)$ Take $G=K_{k,k,\dots,k}$, the complete equipartite graph with $n/k$ partite classes of size $k$.
$(iv)$ If k=n2k=n-2, take $K_{2}+E_{n-2}-2K_{2}$. More generally, for any even $k\leq n-2$, begin with the complete join $K_{n-k}+E_{k-1}$ and supplement it with a further vertex $z$. Delete an edge $vw$ where $v\in K_{n-k}$ and $w\in E_{k-1}$; and insert the new edge $vz$. (For $n=4$, both ways yield $P_{4}$.) $\Box$
Note that k=n1k=n-1 with $k$ even does not admit a connected graph with $\alpha_{\mathrm{od}}=k$, because the only connected graph having $\alpha=n-1$ is the star of degree n1n-1. And, of course, $k=\alpha=n$ implies the union of singleton components, always disconnected.
2.1 Forbidden pair lemma
Here we state a property which is simple but makes many situations concerning odd independent sets more transparent. We introduce two notions.
Definition 7.
(Forbidden pair)  Two nonadjacent vertices $x,y$ form a forbidden pair if they have a common neighbor $z$ such that $N[z]\subseteq N[x]\cup N[y]$.
Definition 8.
(Forcing pair)  Two nonadjacent neighbors $x,y$ of a vertex $z$ form a forcing pair if, for any vertex $w$ in $N(z)$ independent of $x$ and $y$, either $x,w$ or $y,w$ or both are forbidden pairs.
Lemma 9.
If $S\subseteq V(G)$ is an odd independent set, then it contains no forbidden pair, nor a forcing pair.
Proof.
Suppose for a contradiction that $x,y\in S$, where $x,y$ is a forbidden pair, say in regard to a common neighbor $z$, and $S\subseteq V(G)$ is odd independent. By definition, $S$ should contain at least one further neighbor, say $w$, of $z$. However, every neighbor of $z$ belongs to $N[x]\cup N[y]$, hence at least one of $wx\in E(G)$ and $wy\in E(G)$ holds, which is not allowed if $S$ is an independent set.
Assume next that $x,y\in S$ form a forcing pair, say in regard to a common neighbor $z$. Due to odd independence, $S$ has to contain at least one further neighbor, say $w$, of $z$. But then the forbidden pair $x,w$ or $y,w$ is contained in $S$, and we have just proved that this is not possible. $\Box$
Theorem 10.
If $G$ is claw-free, then $\alpha_{\mathrm{od}}(G)=\alpha(G^{2})$ and $\chi_{\mathrm{so}}(G)=\chi(G^{2})$.
Proof.
We prove that a set $S\subseteq V(G)$ is odd independent in $G$ if and only if $S$ is an independent set in $G^{2}$. This will verify both equalities.
The “if” part is clear. For “only if”, consider any independent set $S$ in $G$, and suppose that $S$ is not independent in $G^{2}$. Say, $x,y\in S$ where $xy\notin E(G)$ and $x,y\in N(v)$ for some vertex $v$. Since $G$ is claw-free, the pair $x,y$ dominates all neighbors of $v$. Consequently $N[v]\subseteq N[x]\cup N[y]$. Thus $x,y$ is a forbidden pair, and $S$ cannot be an odd independent set. $\Box$
Remark 11.
The overall presence of forbidden pairs characterizes claw-free graphs. Namely, a graph is claw-free if and only if any two vertices at distance $2$ form a forbidden pair for each of their common neighbors. Indeed, if $G$ contains a claw with center $v$ and leaves $x,y,z$ then $z\in N[v]\,\diagdown\,(N[x]\cup N[y])$, hence $x,y$ is not a forbidden pair.
Further applications of forbidden and forcing pairs are given in [4].
2.2 Applications of $\chi_{\mathrm{so}}(G)$ to lower bounds on $\alpha_{\mathrm{od}}(G)$
In this short subsection we collect the main results from the papers [3, 7, 9, 11] concerning upper bounds on $\chi_{\mathrm{so}}(G)$ and their implications to obtain lower bounds for $\alpha_{\mathrm{od}}(G)$ via the relation $\alpha_{\mathrm{od}}(G)\geq|G|/\chi_{\mathrm{so}}(G)$. This can also serve as a comparison to future results that will be obtained by direct research on $\alpha_{\mathrm{od}}(G)$.
For the next part (A) we use the standard notation $\mathrm{Mad}(G):=\max\,\{d(H)\mid H\subseteq G\}$ for the maximum average degree of $G$.
(A) From the paper [9]
  • Theorem 1.2: If $G$ is a graph with $\mathrm{Mad}(G)\leq 20/7$, then $\chi_{\mathrm{so}}(G)\leq\Delta(G)+4$. $\Longrightarrow$ $\alpha_{\mathrm{od}}(G)\geq n/(\Delta(G)+4)$.
  • Theorem 1.3: If $G$ is a graph with $\Delta(G)\geq 4$ and $\mathrm{Mad}(G)\leq 30/11$, then $\chi_{\mathrm{so}}(G)\leq\Delta(G)+3$. $\Longrightarrow$ $\alpha_{\mathrm{od}}(G)\geq n/(\Delta(G)+3)$.
  • Theorem 1.4: If $G$ is a $C_{4}$-free subcubic graph with $\mathrm{Mad}(G)\leq 30/11$, then $\chi_{\mathrm{so}}(G)\leq 6$. $\Longrightarrow$ $\alpha_{\mathrm{od}}(G)\geq n/6$.
  • Corollary 1.5: For any planar graph $G$, $\chi_{\mathrm{so}}(G)\leq\Delta(G)+4$ if $g(G)\geq 7$, and $\chi_{\mathrm{so}}(G)\leq\Delta(G)+3$ if $g(G)\geq 8$. $\Longrightarrow$ $\alpha_{\mathrm{od}}(G)\geq n/(\Delta(G)+4)$ if $g(G)\geq 7$, and $\alpha_{\mathrm{od}}(G)\geq n/(\Delta(G)+3)$ if $g(G)\geq 8$. These estimates are better than the one in our [3], $n/388$, for $\Delta\leq 384$ or $\leq 385$, respectively.
(B) From the paper [7]
  • Proposition 3: For every outerplanar graph $G$, $\chi_{\mathrm{so}}(G)\leq 8$. $\Longrightarrow$ $\alpha_{\mathrm{od}}(G)\geq n/8$.
  • Theorem 4: For every proper minor-closed graph class $\mathcal{F}$, there exists a constant $c(\mathcal{F})$ such that for every graph $G\in\mathcal{F}$ we have $\chi_{\mathrm{so}}(G)\leq c(\mathcal{F})$. $\Longrightarrow$ $\alpha_{\mathrm{od}}(G)\geq n/c(\mathcal{F})$.
  • Theorem 10: There exists a function $f$ such that for every graph $G$ we have $\chi_{\mathrm{so}}(G)\leq f(\mathrm{tw}(G))$, where $\mathrm{tw}(G)$ denotes the treewidth of $G$. $\Longrightarrow$ $\alpha_{\mathrm{od}}(G)\geq n/f(\mathrm{tw}(G))$.
  • Theorem 19: There exists a function $f$ such that for every graph $G$ we have $\chi_{\mathrm{so}}(G)\leq f(\mathrm{rtw}(G))$, where $\mathrm{rtw}(G)$ denotes the row treewidth of $G$. $\Longrightarrow$ $\alpha_{\mathrm{od}}(G)\geq n/f(\mathrm{rtw}(G))$.
However, the implied constants in Theorems 4, 10, and 19 are not made explicit.
(C) From the paper [11]
  • Generalizing the main results from [7], the following major result is proved.
    Theorem 1.1: Every graph class of bounded expansion has bounded strong odd chromatic number; i.e., for any graph class $\mathcal{F}$ of bounded expansion there is a constant $c(F)$ such that, for every graph $G\in\mathcal{F}$, $\chi_{\mathrm{so}}(G)\leq c(F)$. $\Longrightarrow$ $\alpha_{\mathrm{od}}(G)\geq|G|/c(F)$.
(D) From our paper [3]
  • Proposition 2.1: For every tree $T$ we have $\chi_{\mathrm{so}}(T)\leq 3$, and $\chi_{\mathrm{so}}(T)=2$ holds if and only if $T$ is odd (i.e., if all its vertices have odd degrees). $\Longrightarrow$ $\alpha_{\mathrm{od}}(T)\geq n/3$ and $n/2$, respectively.
  • Proposition 2.2: If $G$ is a connected unicyclic graph other than $C_{5}$, then $\chi_{\mathrm{so}}(G)\leq 4$ $\Longrightarrow$ $\alpha_{\mathrm{od}}(G)\geq n/4$.
  • Corollary 3.5: For every planar graph $G$, $\chi_{\mathrm{so}}(G)\leq 388$. $\Longrightarrow$ $\alpha_{\mathrm{od}}(G)\geq n/388$.
3 Odd independence in some special graphs
In this section we compute, via distinct approaches, the odd independence number for various classes of well-known graphs.
3.1 Paths and cycles
The values $\chi_{\mathrm{so}}(P_{n})$ (equal to 3 for all $n\geq 3$) and $\chi_{\mathrm{so}}(C_{n})$ (at most 4 unless $n=5$) are determined for all $n$ in Section 2 of [3]. The corresponding values of odd independence number are given in the next assertion.
Proposition 12.
We have $\alpha_{\mathrm{od}}(P_{n})=\lceil n/3\rceil$ and $\alpha_{\mathrm{od}}(C_{n})=\lceil(n-2)/3\rceil$.
Proof.
For paths we prove the recursion $\alpha_{\mathrm{od}}(P_{n})=\alpha_{\mathrm{od}}(P_{n-3})+1$ for $n\geq 4$. This, together with the obvious fact $\alpha_{\mathrm{od}}(P_{1})=\alpha_{\mathrm{od}}(P_{2})=\alpha_{\mathrm{od}}(P_{3})=1$, implies $\alpha_{\mathrm{od}}(P_{n})=\lceil n/3\rceil$ by induction.
Let $P_{n}=v_{1}\cdots v_{n}$ be the path of order $n$, and let $S$ be an odd independent set of size $|S|=\alpha_{\mathrm{od}}(P_{n})$ such that the sum $\sum_{v_{i}\in S}i$ is maximum. Consider the vertex $v_{k}\in S$ of highest index in $S$. It must be the case that $k=n$, for otherwise replacing $v_{k}$ with $v_{n}$ we would obtain a set $S^{\prime}$ such that $|S^{\prime}|=|S|$ and also $S^{\prime}$ is odd independent, but the sum of indices in $S^{\prime}$ is larger than that in $S$, contradicting the choice of $S$. Hence $v_{n}\in S$ and $v_{n-1},v_{n-2}\notin S$. Since $\{v_{n-1},v_{n-2}\}$ separates $v_{n}$ from the rest of the path in the strong sense that they have no common neighbors, $S$ is odd independent in $P_{n}$ if and only if so is $S\,\diagdown\,\{v_{n}\}$ in $P_{n-3}=v_{1}\cdots v_{n-3}$. This implies the assertion.
For very short cycles it is clear that $\alpha_{\mathrm{od}}(C_{3})=\alpha_{\mathrm{od}}(C_{4})=\alpha_{\mathrm{od}}(C_{5})=1$. If $n>5$, consider $C_{n}=v_{1}\cdots v_{n}$, and let $S$ be a largest odd independent set. Pick any $v_{i}\in S$. Then $v_{i-1},v_{i-2},v_{i+1},v_{i+2}\notin S$ and $S$ is odd independent in $C_{n}$ if and only if so is $S\,\diagdown\,\{v_{i}\}$ in $C_{n}-\{v_{i-2},v_{i-1},v_{i},v_{i+1},v_{i+2}\}\cong P_{n-5}$ (subscript addition taken modulo $n$). Thus, $\alpha_{\mathrm{od}}(C_{n})=1+\alpha_{\mathrm{od}}(P_{n-5})$ holds, and the assertion follows by the formula on paths. $\Box$
3.2 Moore graphs
Next we consider the famous $d$-regular graphs of diameter 2 for which the Moore bound is tight. It means $C_{5}$ for $d=2$, the Petersen graph for $d=3$, and the Hoffman–Singleton graph [8] for $d=7$. (The existence of a tight construction with $d=57$ on $57^{2}+1=3250$ vertices is still open.) Of course $\alpha_{\mathrm{od}}(C_{5})=1$ and $\chi_{\mathrm{so}}(C_{5})=5$. The other cases are more complicated.
We shall apply the following simple observation.
Lemma 13.
Assume that $G$ has girth at least $5$. If $v\in V(G)$ is any vertex and $S\subseteq N(v)$ any set of odd size, then $S$ is an odd independent set in $G$.
Proof.
By assumption $v$ has an odd number $|S|$ of neighbors in $S$. On the other hand, any $w\in V(G)\,\diagdown\,(S\cup\{v\})$ can have at most one neighbor in $S$, for otherwise it would create a $C_{4}$ together with $v$ and two vertices from $N(v)\cap N(w)$. $\Box$
For simplicity, let us say that an odd independent set $S$ is trivial if $|S|\leq 1$, and non-trivial otherwise.
Proposition 14.
The Petersen graph has $\alpha_{\mathrm{od}}=3$ and $\chi_{\mathrm{so}}=6$. Moreover, every non-trivial odd independent set is the neighborhood $N(v)$ of a vertex $v$.
Proof.
Let $x,y\in S$ be two vertices in any non-trivial odd independent set $S$. Since the diameter is 2, $x$ and $y$ have a common neighbor $w$. Then, by parity reason, $S$ has to contain the third neighbor, say $z$, of $w$. In this case, however, $N(x)\cup N(y)\cup N(z)$ is the entire vertex set, hence no more vertices can occur in $S$. This implies $\alpha_{\mathrm{od}}=3$.
Consider now a strong odd coloring with $\chi_{\mathrm{so}}$ colors. Say, $p$ vertex classes are trivial and $q$ are non-trivial. We then have $p+q=\chi_{\mathrm{so}}$ and $p+3q=10$, implying $q\leq 3$. We observe that the case $q=3$ is impossible. Indeed, $N(v)\cap N(v^{\prime})=\varnothing$ holds for two vertices $v,v^{\prime}$ in the Petersen graph (or more generally in any graph of diameter 2) if and only if $v$ and $v^{\prime}$ are adjacent. Hence, $|S|=3$ would imply the presence of a triangle, while the Petersen graph has girth 5. Consequently $q\leq 2$ and $p+q=10-2q\geq 6$.
A strong odd coloring with 6 colors can be achieved by taking the neighborhoods of two adjacent vertices and four trivial odd independent sets. This completes the proof of $\chi_{\mathrm{so}}=6$. $\Box$
Let us recall that the Hoffman–Singleton graph [8] is 7-regular, has diameter 2 and girth 5, and its order is $7^{2}+1=50$.
Proposition 15.
The Hoffman–Singleton graph has $\alpha_{\mathrm{od}}=15$ and $\chi_{\mathrm{so}}\leq 20$. On the other hand, it has no odd independent set with exactly $k$ vertices in the range $11\leq k\leq 14$.
Proof.
We label the vertices as $0,1,\dots,49$ in the way that $i\mapsto i+10$ (mod 50) is an automorphism, and the first ten vertices have the following adjacencies:
vertex neighbors
0     1, 4, 13, 16, 26, 43, 49
1     0, 2,  6, 18, 28, 36, 47
2     1, 3,  8, 24, 34, 38, 45
3     2, 4, 10, 19, 29, 40, 48
4     0, 3,  5, 22, 32, 35, 46
5     4, 6,  9, 12, 24, 27, 37
6     1, 5,  7, 14, 21, 30, 40
7     6, 8, 11, 19, 25, 35, 49
8     2, 7,  9, 13, 22, 31, 41
9     5, 8, 10, 17, 33, 43, 47
We can generate an odd independent set of size 15 from the triplet $\{0,2,7\}$, taking $S=\{10j,\,10j+2,\,10j+7\mid j=0,1,2,3,4\}$. Every vertex $i\notin S$ has exactly three neighbors [^1] Those 35 neighborhoods form a Kirkman triple system KTS(15)(15) over $S$. in $S$. Similar 15-element odd independent sets are obtained in the same way from the sets $\{0,5,8\}$, $\{1,3,5\}$, $\{1,4,9\}$, $\{2,6,9\}$ also.
Further, we can apply the idea of rotation by 10, that is applying $i\mapsto(i+10)$ mod 50, to construct a strong odd coloring with 20 colors. Let $S_{1}=\{0,2,6,18,47\}$ and $S_{2}=\{1,3,24\}$. These are odd independent sets, being subsets of odd cardinality in $N(1)$ and $N(2)$, respectively. Rotating them with multiples of 10 we obtain 10 mutually disjoint odd independent sets which together cover 40 vertices. The other 10 vertices can be taken as singleton color classes.
Suppose for a contradiction that there exists an odd independent set $S$ of size $k$, $11\leq k\leq 14$. There are $7k$ edges with one end in $S$ and the other in its complement, which we denote by $Q$. In this SSQQ bipartite graph all the q:=50kq:=50-k vertices $i\in Q$ have odd degrees. The most balanced degree distribution of the ends of the $7k\geq 77$ edges among the $50-k\leq 39$ vertices of $Q$ can be determined in two steps. First give 1 to each of the 50k50-k. There remain 8k508k-50 to be distributed, but we must obtain odd degrees everywhere, so the increase cannot be smaller than 2. Hence in the second step we give 2 to $\min\,\{50-k,4k-25\}=4k-25$ vertices of $Q$. This yields 4k254k-25 degrees 3 and (50k)(4k25)=755k(50-k)-(4k-25)=75-5k degrees 1. Each vertex $i\in Q$ having 3 has a neighborhood $N(i)\cap S$ that covers three vertex pairs in $S$; altogether this means at least 12k7512k-75. Observe:
$k$ 11 12 13 14
12k7512k-75 57 69 81 93
$\binom{k}{2}$ 55 66 78 91
Since $12k-75>\binom{k}{2}$ holds in this range, there would occur vertex pairs in $S$ which are contained in the neighborhood of more than one vertex of $Q$. This would imply that the graph contains $K_{2,2}\cong C_{4}$, a contradiction. $\Box$
Also in the lower range, several cardinalities are impossible; for instance, the Hoffman–Singleton graph does not contain any 2-element odd independent set $S$. The reason is that if $u,v\in S$, then $u,v$ are not adjacent, hence they have a common neighbor, say $w$. But then $w$ should have an odd number of further neighbors in $S$.
3.3 Kneser graphs
Theorem 16.
For every $k\geq 2$ and every $n\geq 2k$, the Kneser graph $\mathrm{KG}(n,k)$ has $\alpha_{\mathrm{od}}=\alpha$ if and only if $\binom{n-k-1}{k-1}$ is odd. Equivalently, $\alpha_{\mathrm{od}}=\alpha$ if and only if $\binom{n-1}{k-1}\not\equiv\sum_{t=1}^{k}\binom{k}{t}\binom{n-k-1}{k-1-t}$ (mod 2)($\mathrm${mod}\ 2).
Proof.
We first note that $\mathrm{KG}(2k,k)$ is a matching, i.e. regular of degree 1. Thus, every independent set in $\mathrm{KG}(2k,k)$ is also odd independent and $\alpha_{\mathrm{od}}=\alpha$ holds.
Assume $n>2k$. A largest independent set in $\mathrm{KG}(n,k)$ corresponds to a collection of the maximum number of pairwise intersecting $k$-element subsets of an $n$-element set. According to the famous Erdős–Ko–Rado theorem, all such largest families $\mathcal{F}$ are obtained by picking all $k$-sets containing a fixed element. Every $k$-element set $F\notin\mathcal{F}$ is disjoint from exactly $\binom{n-k-1}{k-1}$ members of $\mathcal{F}$, and this is exactly the degree of vertices in $\mathrm{KG}(n,k)$. This is odd if and only if the largest independent sets in $\mathrm{KG}(n,k)$ are also odd independent.
The second condition is equivalent to the first one because the equality $\sum_{t=1}^{k}\binom{k}{t}\binom{n-k-1}{k-1-t}=\binom{n-1}{k-1}-\binom{n-k-1}{k-1}$ is valid for all $n\geq 2k$. $\Box$
3.4 The graphs $S(K_{n})$, $H_{n,n}$, $K_{n}\,\Box\,K_{n}$
$S(K_{n})$ — the complete subdivision of $K_{n}$.
For $n\geq 2$ the graph $S(K_{n})$ is obtained from $K_{n}$ by subdividing each edge with a new “subdivision vertex”. Hence it has $\binom{n}{2}+n=\binom{n+1}{2}$ vertices, $n$ of which have degree n1n-1 and $\binom{n}{2}$ have degree 2. The $n$ vertices from $K_{n}$ will be referred to as “original vertices”.
Proposition 17.
If $n\geq 2$ is any integer, then:
  • $(i)$
    $\chi_{\mathrm{so}}(S(K_{n}))=n$ for all $n$, except that $\chi_{\mathrm{so}}(S(K_{2}))=3$ and $\chi_{\mathrm{so}}(S(K_{4}))=5$.
  • $(ii)$
    $\alpha_{\mathrm{od}}(S(K_{n}))=\binom{n}{2}$ if $n$ is even, and $\alpha_{\mathrm{od}}(S(K_{n}))=\binom{n-1}{2}+1$ if $n$ is odd.
Proof.
Denote the original vertices by $v_{0},\dots,v_{n-1}$, and let $x_{i,j}$ be the subdivision vertex corresponding to the edge $v_{i}v_{j}$ for every $0\leq i<j\leq n-1$. Observe that no odd independent set $S$ can contain more than one original vertex, because if $v_{i},v_{j}\in S$ held, then $x_{i,j}$ would have exactly two neighbors in $S$. This property immediately implies $\chi_{\mathrm{so}}(S(K_{n}))\geq n$, and also $\alpha_{\mathrm{od}}(S(K_{n}))\leq\binom{n}{2}$ because a larger $S$ would require the simultaneous presence of all subdivision vertices together with one original vertex.
Proof of $(i)$ for odd $n$. We label the color classes $S_{i}$ with $0,1,\dots,n-1\in\mathbb{Z}_{n}$. Define $S_{i}=\{v_{i}\}\cup\{x_{i+j,i-j}\mid j\in\mathbb{Z}_{n}\,\diagdown\,\{0\}\}$, subscript addition taken modulo $n$. Then the closed neighborhood of each original vertex contains each color from $0,1,\dots,n-1$ exactly once, and the two neighbors of each subdivision vertex have distinct colors, hence a strong odd coloring with $n$ colors is obtained.
Proof of $(i)$ for even $n$. We begin with noting that a strong odd coloring with $n$ colors is not possible for $n=2$ and $n=4$. Indeed, $S(K_{2})\cong P_{3}$, which needs exactly three colors. Moreover, assigning color $i$ to the original vertex $v_{i}$, we cannot use this color for any $x_{i,j}$ ($j\neq i$), so the possible choices of subdivision vertices of color $i$ correspond to a triangle in $K_{4}$, and at most one such appearance is possible (as any two triangles of $K_{4}$ share an edge). Then the odd-degree condition with respect to $S$ implies that only one $x_{i,j}$ can get color $i$. This means that at least two of the six subdivision vertices must have colors distinct from $0,1,2,3$. On the other hand, this is feasible with five colors, e.g. assign color $i$ to $v_{i}$ and $x_{i+1,i+2}$ ($i=0,1,2,3$) and color 4 to $x_{0,2}$ and $x_{1,3}$.
Assume that $n\geq 6$. We prove that $n$ colors suffice. Also in this case we assign color $i$ ($i=0,1,\dots,n-1$) to the original vertex $v_{i}$. The subdivision vertices are colored as follows:
  • $x_{i,j}$ has color n1n-1 for all $0\leq i<j\leq n-3$.
  • $x_{i,n-2}$ has color 0 for all $1\leq i\leq n-3$.
  • $x_{i,n-1}$ has color n2n-2 for all $0\leq i\leq n-4$.
  • $x_{0,n-2}$ and $x_{n-3,n-1}$ have color 1.
  • $x_{n-2,n-1}$ has color n3n-3.
This coloring corresponds to an edge decomposition of $K_{n}$ in which the edge classes are isomorphic to $K_{n-2}$, $K_{1,n-3}$, $K_{1,n-3}$, $2K_{2}$, and $K_{2}$, respectively. Hence a strong odd coloring of $S(K_{n})$ is obtained.
Proof of $(ii)$. If $n$ is even, then every original vertex is adjacent to an odd number of subdivision vertices, hence $\{x_{i,j}\mid 0\leq i<j\leq n-1\}$ is an odd independent set of size $\binom{n}{2}$. If $n$ is odd, we take the odd independent set $\{v_{0}\}\cup\{x_{i,j}\mid 1\leq i<j\leq n-1\}$ of size $\binom{n-1}{2}+1$, and claim that no larger one exists.
Let $E(S)$ denote the set of edges in $K_{n}$ corresponding to the subdivision vertices belonging to $S$. If $S$ is odd independent, then all degrees in the graph induced by $E(S)$ are odd, hence this graph must have an even order. This means at most n1n-1 because $n$ is odd. Thus, $|E(S)|\leq\binom{n-1}{2}$. In addition, $S$ can contain at most one original vertex, as told at the beginning of the proof. This implies the validity of the assertion. $\Box$
This construction yields that, contrary to 1-degenerate graphs, and also in sharp contrast with the classical chromatic number, $\chi_{\mathrm{so}}$ is unbounded in the class of 2-degenerate graphs.
Corollary 18.
There exist an infinite sequence of $2$-degenerate graphs $G$ such that $\chi_{\mathrm{so}}(G)$ grows proportionally to $\sqrt{|G|}$.
$H_{n,n}$ — the Half-graph.
The graphs $S(K_{n})$ are examples of bipartite graphs with large $\alpha_{\mathrm{od}}$ and moderate $\chi_{\mathrm{so}}$.
The counterpart pathology is the half-graph $H_{n,n}$ with vertex set $A\cup B$, where $A=\{u_{1},\dots,u_{n}\}$, $B=\{v_{1},\dots,v_{n}\}$; vertex $u_{i}$ is adjacent to vertex $v_{j}$ for all $j\geq i$.
Proposition 19.
$\chi_{\mathrm{so}}(H_{n,n})=n+1$ and $\alpha_{\mathrm{od}}(H_{n,n})=2$, all odd independent sets having the form $\{v_{i},u_{j}\}$, $i<j$.
Proof.
Suppose for a contradiction that an odd independent set $S$ has more than one vertex in $B$. Let $i<j$ be its two highest indices in $B$. Then $u_{i}\in A$ is adjacent just to $v_{i}$ and $v_{j}$, contradicting the conditions on odd independent sets. Hence $|S\cap B|=1$, and in the same way looking at the hypothetical two lowest indices in $A$ we also obtain $|S\cap A|=1$. Consequently every odd independent set larger than a singleton is a nonadjacent vertex pair $v_{i},u_{j}$, $i<j$.
Concerning any strong odd coloring this implies that it is a color partition into an AABB matching of non-edges, together with singletons. Since $u_{1}$ and $v_{n}$ have degree $n$, at most n1n-1 non-singleton classes can occur, hence $\chi_{\mathrm{so}}(H_{n,n})>n$. On the other hand, $n+1$ is achievable by taking $u_{1}$ and $v_{n}$ as singletons, with the further odd independent sets $\{u_{i},v_{i-1}\}$ for $i=2,3,\dots,n$. $\Box$
$K_{n}\,\Box\,K_{n}$ — Cartesian product of complete graphs.
This graph is opposed to $S(G)$, which has been seen to have large $\alpha_{\mathrm{od}}$ and moderate $\chi_{\mathrm{so}}$.
Proposition 20.
For $H=K_{n}\,\Box\,K_{n}$, $\alpha_{\mathrm{od}}(H)=1$ and $\chi_{\mathrm{so}}(H)=|H|\sim(\Delta(H))^{2}\!/4$.
Proof.
If $u,v\in H$ are nonadjacent vertices, then they have precisely two neighbors, say one of them is $w$. However, $N(w)$ induces $2K_{n-1}$ in $H$, consequently $\{u,v\}$ cannot be extended to any independent set containing three vertices from $N(w)$. This implies $\alpha_{\mathrm{od}}(H)=1$. Since $\alpha_{\mathrm{od}}(G)\chi_{\mathrm{so}}(G)\geq|G|$ holds for every $G$, putting $G=H$ and using the trivial inequality $\chi_{\mathrm{so}}(H)\leq|H|$ we also obtain $\chi_{\mathrm{so}}(H)=|H|$. The asymptotic formula follows as $H$ is (2n2)(2n-2)-regular. $\Box$
One can observe that the same proof gives $\alpha_{\mathrm{od}}(H)=1$ and $\chi_{\mathrm{so}}(H)=|H|$ more generally for $H=K_{p}\,\Box\,K_{q}$. Note further that $G=K_{p}\,\Box\,K_{q}$ is claw-free and $G^{2}\cong K_{pq}$, hence Proposition 20 also follows directly from Theorem 10.
4 Automorphisms and recursive lower bound
In this section we present a recursive method to obtain lower bounds on $\alpha_{\mathrm{od}}$ for graphs satisfying a certain symmetry property. An application of this approach to cube graphs will be given.
For this purpose we define a binary operation $\mu$ acting on pairs of graphs. Let $G$ be a vertex-labeled graph and $H$ any graph. Define $\mu[G,H]$ as the graph obtained as follows. Replace each vertex of $w\in V(H)$ with a copy $G_{w}=(V_{w},E_{w})$ of $G$, and between those copies which are adjacent in $H$, take a perfect matching that connects vertices of the same label.
Theorem 21.
Let $G$ and $H$ be graphs with the following properties:
  • $G$ admits an automorphism $\eta$ such that $\eta(v)\in N(v)$ for every $v\in V(G)$, and the length of each orbit of $\eta$ is even;
  • $H$ is bipartite and each $w\in V(H)$ has an even degree.
Then $\alpha_{\mathrm{od}}(\mu[G,H])\geq|H|\cdot\alpha_{\mathrm{od}}(G)$.
Proof.
If $S$ is an odd independent set in $G$, then by automorphism so is $S^{*}:=\eta(S)$ as well. Let $A,B$ be the two vertex classes of $H$, and consider an $S$ with $|S|=\alpha_{\mathrm{od}}(G)$. If $w\in A$, let $S_{w}$ be the copy of $S$ in $G_{w}$; and if $w\in B$, let $S_{w}$ be the copy of $S^{*}$ in $G_{w}$. We select $S^{+}:=\cup_{w\in V(H)}S_{w}$. Clearly, $|S^{+}|=|H|\cdot\alpha_{\mathrm{od}}(G)$. The proof will be done if we prove that $S^{+}$ is an odd independent set.
Trying to locate a hypothetical edge $u_{1}u_{2}$ inside $S^{+}$, there should exist $w_{1},w_{2}$ such that $u_{i}\in S_{w_{i}}$ ($i=1,2$, $w_{1}\neq w_{2}$) and $w_{1}w_{2}\in E(H)$, because each $S_{w}$ is independent and there is no $G_{w_{1}}$– $G_{w_{2}}$ edge for $w_{1}w_{2}\notin E(H)$. But then $u_{1}u_{2}$ should connect two copies of a $u\in V(G)$, which cannot be the case because one of them is in $A$ while the other is in $B$, hence $u_{1}\mapsfrom u$ and $u_{2}\mapsfrom\mu(u)\in N(u)$ are nonadjacent in $\mu[G,H]$.
It remains to show that every vertex of $V(\mu[G,H])\,\diagdown\,S^{+}$ adjacent to $S^{+}$ has an odd number of neighbors in $S^{+}$. By symmetry reason, it suffices to consider vertices in the copy $G_{w_{1}}$ of $G$. Assume that $v_{1}\in(G_{w_{1}})\,\diagdown\,S_{w_{1}}$ is the copy of $v\in V(G)\,\diagdown\,S$.
If $v$ has no neighbor in $S$, then referring to the inverse automorphism $\mu^{-1}$ we see that $\mu(v)$ has no neighbor in $S^{*}$, thus $v_{1}$ is not adjacent to $S^{+}$ either. The other possibility is that $v$ has an odd number of neighbors in $S$. This yields that $v_{1}$ has an odd number of neighbors in $S_{w_{1}}$. All the other neighbors $u_{i}\in S^{+}$ of $v_{1}$, i.e. those in any $S_{w_{i}}$ where $w_{i}\in B$, are copies of the same vertex $u\in S^{*}$ where $u=\mu(v)$ holds by the definition of $\mu[G,H]$. Since all degrees in $H$ are even, we obtain $|N(v_{1})\cap S^{+}|\equiv|N(v_{1})\cap S_{w_{1}}|$ (mod 2), hence odd. It follows that $S^{+}$ is an odd independent set. $\Box$
In a similar way we can also obtain:
Theorem 22.
For any graph $G$, if $H$ is a bipartite graph with all vertex degrees odd, then $\alpha_{\mathrm{od}}(\mu[G,H])\geq\frac{1}{2}|H|\cdot\alpha_{\mathrm{od}}(G\,\Box\,K_{2})$.
Proof.
We apply notation and several ideas from the previous proof. Let $A$ and $B$ be the two vertex classes of $H$, and $S\cup S^{*}$ an odd independent set of maximum size $|S|+|S^{*}|=\alpha_{\mathrm{od}}(G\,\Box\,K_{2})$ in $G\,\Box\,K_{2}$, where $S$ is the subset in one copy of $G$ and $S^{*}$ is the subset in the other copy. We may assume $|A|\geq|B|$ and $|S|\geq|S^{*}|$ without loss of generality.
As in the previous proof, inside the copies $G_{w}\subset\mu[G,H]$ we select the copies of $S$ or $S^{*}$ if $w\in A$ or $w\in B$, respectively. The argument in the previous proof clearly yields that the obtained set $S^{+}$ is independent. We need to prove that it is also odd independent. Let $w_{1}w_{2}\in E(H)$ be any edge, $w_{1}\in A$, and let $v\in V(G)\,\diagdown\,S$ be any vertex with at least one neighbor in $S\cup S^{*}$. Consider the copy $v_{1}$ of $v$ in $G_{w_{1}}$. The number of its neighbors in $V(G_{w_{1}})\cup V(G_{w_{2}})$ is odd, because the subgraph induced by this union is isomorphic to $G\,\Box\,K_{2}$ and $S\cup S^{*}$ is odd independent. The additional $\deg_{H}(w)-1$ neighbors of $w_{1}$ increase this with $(\deg_{H}(w)-1)\cdot|N(v)\cap S^{*}|$, hence adding it to $|N(v)\cap(S\cup S^{*})|$ the sum remains odd, due to the degree assumption on $H$. The same is true for the copy $v_{2}$ of $v$ in $G_{w_{2}}$. Thus, $S^{+}$ is an odd independent set.
We now have $\alpha_{\mathrm{od}}(\mu[G,H])\geq|S^{+}|=|A|\cdot|S|+|B|\cdot|S^{*}|=\frac{1}{2}(|A|+|B|)|S|+\frac{1}{2}(|A|-|B|)|S|+|B|\cdot|S^{*}|\geq\frac{1}{2}(|A|+|B|)|S|+\frac{1}{2}(|A|-|B|)|S^{*}|+|B|\cdot|S^{*}|=\frac{1}{2}(|A|+|B|)(|S|+|S^{*}|)=\frac{1}{2}|H|\cdot\alpha_{\mathrm{od}}(G\,\Box\,K_{2})$. $\Box$
Remark 23.
Although the two lower bounds above have been derived using very similar approach, the estimates are independent in the qualitative sense, because $\alpha_{\mathrm{od}}(G\,\Box\,K_{2})$ is not expressible as a function of $\alpha_{\mathrm{od}}(G)$. This fact is demonstrated by the following simple examples:
  • $\alpha_{\mathrm{od}}(K_{2})=1$ and $\alpha_{\mathrm{od}}(K_{2}\,\Box\,K_{2})=1$. ($K_{2}\,\Box\,K_{2}\cong C_{4}$.)
  • $\alpha_{\mathrm{od}}(C_{4})=1$ and $\alpha_{\mathrm{od}}(C_{4}\,\Box\,K_{2})=4$. ($C_{4}\,\Box\,K_{2}\cong Q_{3}$ is 3-regular bipartite.)
We can offer a further approach based on the following notion. Due to its generality there may occur interesting applications of it, although we have not found any so far.
Definition 24.
Let $G$ be a graph and $B$ an odd independent set in $G$. We say that $B$ is odd-suitable if there is a partition $V(G)=V_{1}\cup V_{2}$ such that the sets $B_{1}:=B\cap V_{1}$, $B_{2}:=B\cap V_{2}$ and $B_{1}^{*}:=N(B_{1})\cap V_{2}$, $B_{2}^{*}:=N(B_{2})\cap V_{1}$ satisfy the following conditions for $i=1,2$ :
  • $N(B_{i}^{*})\cap V_{i}=B_{i}$,
  • the $B_{i}$$B_{i}^{*}$ edges induce an odd bipartite graph.
In the above theorems concerning $\alpha_{\mathrm{od}}(\mu[G,H])$ the $B_{i}$$B_{i}^{*}$ subgraphs are just matchings. In this way the notion of odd-suitable sets opens the road to further control on $\alpha_{\mathrm{od}}$ in certain graph classes built recursively. Currently we do not have applications of this kind, therefore we leave it for future research to develop the theory in this direction.
5 Application to hypercubes
Let $Q_{d}$ denote the $d$-dimensional hypercube, whose vertex set is $\{0,1\}^{d}$, and two vertices are adjacent if and only if they differ in exactly one coordinate. Equivalently, one can take the subsets of $\{1,\dots,d\}$ as vertices, two subsets being adjacent if and only if their symmetric difference is a single element. In this representation the mapping $\eta$ with $F\leftrightarrow F\cup\{d\}$ for all $F\subseteq\{1,\dots,d-1\}$ satisfies the conditions of Theorem 21, all orbits having length 2. Further, one may also view $Q_{d}$ as $\mu[Q_{d-1},K_{2}]\cong Q_{d-1}\Box K_{2}$ or $Q_{d-2}\,\Box\,C_{4}\cong\mu[Q_{d-2},C_{4}]$ (cf. Theorem 22 with 1-regular $K_{2}$ and Theorem 21 with 2-regular $C_{4}$).
Before turning to odd independent sets, let us note that the strong odd chromatic number of hypercubes is easy to determine as follows.
Proposition 25.
$\chi_{\mathrm{so}}(Q_{d})=2$ if $d$ is odd, and $\chi_{\mathrm{so}}(Q_{d})=4$ if $d$ is even.
Proof.
If $d$ is odd, then the unique proper vertex 2-coloring of $Q_{d}$ is also a strong odd coloring, hence the first formula is clear. If $d$ is even, then $Q_{d}$ decomposes to two copies of $Q_{d-1}$ with a perfect matching betrween them. Hence a proper coloring with colors $1,2$ in the first copy and with colors $3,4$ in the second copy is a strong odd coloring with four colors. Fewer colors are not enough, because the odd chromatic number of $Q_{d}$ for even $d$ is equal to 4, as proved in [3, Theorem 3.2], and $\chi_{\mathrm{so}}(G)\geq\chi_{\mathrm{o}}(G)$ holds for all graphs $G$. $\Box$
From Proposition 3 we obtain:
Corollary 26.
If $d$ is odd, then $\alpha_{\mathrm{od}}(Q_{d})=\alpha(Q_{d})=2^{d-1}$. If $d$ is even, then $\alpha_{\mathrm{od}}(Q_{d})\leq\bigl(1-\frac{1}{2d-1}\bigr)\cdot 2^{d-1}$.
We do not know, however, whether or not the upper bound for the case of even $d$ is tight. We can prove some partial results.
Proposition 27.
The following lower bounds and equalities are valid.
  • $(i)$
    $\alpha_{\mathrm{od}}(Q_{4k})\geq 2\sum_{i=1}^{k}\binom{4k-1}{2i-1}$ for all $k\geq 1$.
  • $(ii)$
    $\alpha_{\mathrm{od}}(Q_{4})=6$, $\alpha_{\mathrm{od}}(Q_{6})=24$, $\alpha_{\mathrm{od}}(Q_{8})=112$.
  • $(iii)$
    $\alpha_{\mathrm{od}}(Q_{d})\geq\frac{7}{16}\cdot 2^{d}$ for all even $d\geq 8$.
Proof.
We represent the vertices of $Q_{d}$ as subsets of $\{1,\dots,d\}$.
$(i)$ To obtain an odd independent set achieving the claimed size, we select those subsets of size $1,3,\dots,2k-1$ which do not contain the element $4k$, and those of size $2k+1,2k+3,\dots,4k-1$ which contain $4k$.
If a $2k$-element set does not contain $4k$, it has $2k$ neighbors of size 2k12k-1 (omitting any one of its elements), and a unique neighbor of size $2k+1$ (inserting $4k$ into it); together these are $2k+1$ selected neighbors. If a $2k$-element set contains $4k$, then it has 2k12k-1 neighbors of size 2k12k-1 (omitting any one element other than $4k$), and $2k$ neighbors of size $2k+1$ (inserting any one element from its complement); together these mean 4k14k-1 selected neighbors in $Q_{4k}$. Small sets of even size $2s<2k$ do not contain $4k$, hence they have $2s$ neighbors of size 2s12s-1 and 4k2s14k-2s-1 neighbors of size $2s+1$; together these are 4k14k-1. Large even sets of size 4k2s4k-2s are complementary to small ones, so they also have 4k14k-1 selected neighbors. Thus, an odd independent set is obtained, which has the claimed number of elements.
$(ii)$ The lower bound 6 is the first case of $(i)$, and 24 is obtained by the $\mu[G,H]$ construction with $H=C_{4}$. The lower bound 112 was obtained via computer aid. Also tightness has been verified by a computer code. The binary sequences describing the characteristic vectors of the 112 sets are as follows:
00000010, 00000100, 00000111, 00001000, 00001011, 00001101, 00001110,
00010000, 00010011, 00010101, 00010110, 00011001, 00011010, 00011100,
00100000, 00100011, 00100101, 00100110, 00101001, 00101100, 00101111,
00110001, 00110010, 00110111, 00111000, 00111011, 00111101, 00111110,
01000000, 01000011, 01000101, 01001001, 01001010, 01001100, 01001111,
01010001, 01010010, 01010100, 01010111, 01011011, 01011101, 01011110,
01100001, 01100010, 01100100, 01100111, 01101000, 01101011, 01101110,
01110000, 01110101, 01110110, 01111001, 01111010, 01111100, 01111111,
10000000, 10000011, 10000101, 10000110, 10001001, 10001010, 10001111,
10010001, 10010100, 10010111, 10011000, 10011011, 10011101, 10011110,
10100001, 10100010, 10100100, 10101000, 10101011, 10101101, 10101110,
10110000, 10110011, 10110101, 10110110, 10111010, 10111100, 10111111,
11000001, 11000010, 11000100, 11000111, 11001000, 11001101, 11001110,
11010000, 11010011, 11010110, 11011001, 11011010, 11011100, 11011111,
11100011, 11100101, 11100110, 11101001, 11101010, 11101100, 11101111,
11110001, 11110010, 11110100, 11110111, 11111000, 11111011, 11111101.
A lower bound 104 for $d=8$ can be obtained by the following construction, that has a transparent structure. We take $\{1,2,...,8\}$ as the underlying set whose $2^{8}$ subsets are the vertices of $Q_{8}$. Select all 4-element subsets, take those 2-element subsets which form the 4-regular bipartite Turán graph (16 pairs), the 16 corresponding 6-element sets whose complements form the Turán graph, and finally the empty set and the entire set $\{1,2,...,8\}$. These are $70+16+16+1+1=104$ vertices in $Q_{8}$.
A triplet has 5 extensions to a selected quadruplet, and contains either 2 or 0 selected pairs, depending on whether it intersects both $\{1,2,3,4\}$ and $\{5,6,7,8\}$ or not. Hence, in $Q_{8}$ the number of selected neighbors of a triplet is 7 or 5. A singleton has 4 extensions to a selected pair and contains the empty set, hence its number of selected neighbors is 5. The degrees of 5-sets and 7-sets are analogous. Thus, the selected sets generate an odd independent set in $Q_{8}$.
$(iii)$ This estimate follows by Theorem 21 with $H=C_{4}$. $\Box$
6 Complements of triangle-free graphs
Here we investigate graphs $G$ whose complement $\overline{G}$ is triangle-free. We will analyze the possible combinations of the diameter of such a graph and its complement.
Let us begin with the following two complexity observations.
Proposition 28.
If $\overline{G}$ is triangle-free, then $\chi_{\mathrm{so}}(G)$ can be determined in polynomial time.
Proof.
In any proper coloring (not only in strong odd, but also in non-restricted) of $G$, every color class is either a singleton or a vertex-pair, because $\alpha(G)=2$. Hence we can detect all odd-independent vertex pairs by brute force, their number is $O(n^{2})$. These pairs will be the edges of an auxiliary graph $H$. This $H$ has matching number $m$ if and only if $G$ has $\chi_{\mathrm{so}}(G)=n-m$, where $2m$ vertices are covered by the $m$ matching edges and the remaining n2mn-2m vertices form singleton color classes. So, we can construct $H$ in polynomial time and then determine its matching number efficiently, applying some of the well-known algorithms. $\Box$
Proposition 29.
In the family $\mathcal{F}_{k}$ of all graphs with $\alpha(G)\leq k$, determining $\alpha_{\mathrm{od}}(G)$ and $\alpha(G^{2})$ can be done in polynomial time for every fixed $k$.
Proof.
This can be done by checking all $j$-vertex subsets, $j=1,\dots,k$, (a) for independence, (b) for odd independence, and (c) for independence in $G^{2}$. $\Box$
The family $\mathcal{F}_{2}$ is the family of the complements of triangle-free graphs, as a graph $G$ has $\alpha(G)\leq 2$ if and only if its complement $\overline{G}$ is triangle-free. So, in $\mathcal{F}_{2}$ we expect that some structural results (beyond the immediate complexity result) are also possible.
We now turn to the analysis of $\mathrm{diam}(G)$ and $\mathrm{diam}(\overline{G})$, and their relation to odd independence. Under the triangle-free assumption on $G$ we are going to prove:
  • If $\mathrm{diam}(G)\geq 4$, then $\alpha_{\mathrm{od}}(\overline{G})=\alpha((\overline{G})^{2})=1$ and $\chi_{\mathrm{so}}(G)=|G|$.
  • If $\mathrm{diam}(G)=3$ and $\mathrm{diam}(\overline{G})=2$, then $\alpha_{\mathrm{od}}(\overline{G})=\alpha((\overline{G})^{2})=1$ and $\chi_{\mathrm{so}}(\overline{G})=|G|$.
  • If $\mathrm{diam}(G)=3$ and $\mathrm{diam}(\overline{G})=3$, then $\alpha(\overline{G})=\alpha_{\mathrm{od}}(\overline{G})=\alpha((\overline{G})^{2})=2$.
  • If $\mathrm{diam}(G)=2$ and $\mathrm{diam}(\overline{G})=2$, then $\alpha_{\mathrm{od}}(\overline{G})=\alpha((\overline{G})^{2})=1$ and $\chi_{\mathrm{so}}(\overline{G})=|G|$.
  • (1)
    $\mathrm{diam}(G)\geq 4$, $G$ triangle-free
In this case it is well known that $\mathrm{diam}(\overline{G})\leq 2$; see, e.g., [2]. However, then in $\overline{G}$, no odd independent set can have two vertices because the diameter is at most 2 and, as the independence number is at most 2, we conclude that $\alpha_{\mathrm{od}}(\overline{G})=\alpha((\overline{G})^{2})=1$ and $\chi_{\mathrm{so}}(\overline{G})=|G|$.
Hence this graph class provides another example for $\alpha_{\mathrm{od}}(G)=1$ or, equivalently, for $\chi_{\mathrm{so}}(G)=|G|$.
A large class of further examples will be obtained below at the end of part (3).
  • (2)
    $\mathrm{diam}(G)=3$, $\mathrm{diam}(\overline{G})=2$, $G$ triangle-free
Then $\alpha_{\mathrm{od}}(\overline{G})=\alpha(\overline{G}^{2})=1$ and $\chi_{\mathrm{so}}(\overline{G})=|G|$.
  • (3)
    $\mathrm{diam}(G)=3$, $\mathrm{diam}(\overline{G})=3$, $G$ triangle-free
For such graphs, two vertices realizing the diameter of $\overline{G}$ form both a maximum independent set and also an odd independent set. Further, $\alpha(\overline{G})=\alpha_{\mathrm{od}}(\overline{G})=\alpha(\overline{G}^{2})=2$.
Furthermore, if $\mathrm{diam}(G)=3$, then we know $\mathrm{diam}(\overline{G})\leq 3$. Below we give constructions for both $\mathrm{diam}(\overline{G})=3$ and $\mathrm{diam}(\overline{G})=2$.
Construction of triangle-free graphs with diameter 3 whose complements have diameter 3.
Such graphs exists, the smallest one is the self-complementary $G=P_{4}$ with $\alpha_{\mathrm{od}}(G)=\alpha(G)=2$, cf. Problem 33 $(i)$. Also the graph $G=C_{6}=v_{1}v_{2}\dots v_{6}$ supplemented with the edges $(v_{1},v_{4})$ and $(v_{2},v_{5})$ is triangle-free of diameter 3. Its complement consists of two triangles connected by an edge, hence having diameter 3 and $\alpha_{\mathrm{od}}(\overline{G})=\alpha(\overline{G})=2$ (but $\alpha(G)=3$).
More generally, consider $H=K_{n,m}-e$ (one edge omitted), with $n\geq m\geq 2$. This $H$ is triangle-free and has $\mathrm{diam}(H)=3$, forced by the vertices of the removed edge. Its complement $G=\overline{H}$ is $K_{n}\cup K_{m}$ supplemented with one edge connecting the two complete parts, hence having diameter 3. Further, $\alpha_{\mathrm{od}}(G)=\alpha(G)=2$ (cf. again Problem 33).
This construction can be extended further, taking $K_{n,m}$ ($n\geq m\geq 2$) and deleting $tK_{2}$ where $1\leq t\leq m-1$. This is a triangle-free graph (in fact bipartite) of diameter 3. The complement is the union of $K_{n}$ and $K_{m}$ connected by $tK_{2}$ between them. Hence there are vertices $u$ in $K_{n}$ and $v$ in $K_{m}$ where $v$ is not connected directly by any edge to $K_{n}$ while $u$ is not connected directly by any edge to $K_{m}$. Then the distance between $u$ and $v$ is 3. We take $u$ and $v$ as an odd independent set, which is also obviously a maximum independent set, establishing $\alpha_{\mathrm{od}}=\alpha$.
Construction of triangle-free graphs with diameter 3 whose complements have diameter 2.
In this case $\alpha_{\mathrm{od}}(\overline{G})=1$ and $\chi_{\mathrm{so}}(\overline{G})=|G|$ holds (cf. Problem 35).
Let $G$ be a triangle-free graph with $\mathrm{diam}(G)=2$, $|G|\geq 3$, and $\Delta(G)\leq|G|-2$. We construct $H=G\,\Box\,K_{2}$, i.e., two vertex-disjoint copies of $G$ supplemented with a perfect matching between the corresponding vertices. This $H$ is a triangle-free graph with diameter 3.
We claim that the complement $\overline{H}$ of $H$ has diameter 2. In this graph the two copies of $\overline{G}$ are joined by a complete bipartite graph from which just a matching has been removed. So, inside each $\overline{G}$, any two vertices have a common neighbor in the opposite copy of $\overline{G}$, because $|G|\geq 3$.
Consider two vertices $u,v$ in distinct copies of $\overline{G}$. If they are copies of distinct vertices of $G$, then they are directly adjacent in $\overline{H}$. Otherwise, if they are copies of the same vertex say $w$, we know that $\deg_{G}(w)\leq|G|-2$, hence $u$ has a neighbor $z$ in its $\overline{G}$ and then uzvuzv is an uuvv path of length 2. Thus, $\mathrm{diam}(\overline{H})=2$ holds, indeed.
Observe that this method can be generalized to constructions of triangle-free graphs of any diameter $t\geq 3$ whose complements have diameter 2. Just start with a triangle-free graph $G$ having $\mathrm{diam}(G)=t$ and take $H=G\,\Box\,K_{2}$. The above argument verifies $\mathrm{diam}(\overline{H})=2$, hence covering case (1) in a general way.
  • (4)
    $\mathrm{diam}(G)=2$, $G$ triangle-free
The smallest such graph is $C_{5}$. We will prove that under the assumptions of (4) either $\mathrm{diam}(\overline{G})=2$ or $\overline{G}$ is disconnected. So, in the following we exclude finite diameter greater than 2.
  • (4.a)
    $\mathrm{diam}(\overline{G})\geq 4$
If $\overline{G}$ is connected and has diameter at least 4, then there are vertices $v_{1},v_{2},v_{3},v_{4},v_{5}$ forming an induced path. But then $\{v_{1},v_{3},v_{5}\}$ is an independent set in $\overline{G}$, contradicting the assumption that $G$ is triangle-free.
  • (4.b)
    $\mathrm{diam}(\overline{G})=3$
Suppose that the induced path xuvyxuvy realizes diameter 3 in $\overline{G}$. If $x$ has another neighbor $z$, then $z$ must be adjacent to $u$ as otherwise $\{z,u,y\}$ is an independent set of cardinality 3 in $\overline{G}$, which is forbidden. Similarly, if $z$ is a neighbor of $y$, then it is also adjacent to $v$. For the same reason both $N[x]$ and $N[y]$ induce complete subgraphs in $\overline{G}$. What is more, the union of these two cliques is the entire vertex set, for otherwise a vertex $z$ not in their union would form a 3-element independent set with $x$ and $y$. Thus, $\overline{G}\supseteq(K_{a}\cup K_{b})+e$ where $e$ is the edge $uv$ and $a+b=|G|$ (namely, $a=\deg(x)+1$, $b=\deg(y)+1$). But then $G\subseteq K_{a,b}-e$ holds, which implies $\mathrm{diam}(G)\geq 3$, hence contradicting the basic assumption of (4).
Construction of triangle-free graphs with diameter 2 whose complements have diameter 2.
Start with $K_{n,m}$, $n\geq m\geq 2$, fix a matching of $t$ edges, $1\leq t\leq m$, and subdivide these $t$ matching-edges into paths of length 2; call this graph $G=G(n,m,t)$. This is a triangle-free graph of diameter 2. (Note that $m>1$ is mandatory, otherwise the graph would have diameter 3.)
The complement $H=H(n,m,t)$ of $G(n,m,t)$ is connected, thanks to $m>1$ and to the middle vertex of any subdivided edge. Hence, by the analysis above, we know without further checking that $\mathrm{diam}(\overline{G})=2$.
7 Graphs with $\chi_{\mathrm{so}}(G)=\Delta^{2}(G)+1$ and $\alpha_{\mathrm{od}}(G)\geq|G|/(\Delta^{2}(G)+1)$
For a simpler notation, we will write $\Delta^{2}(G)$ for $(\Delta(G))^{2}$ throughout.
We have already seen that $\chi_{\mathrm{so}}(G)\leq\chi(G^{2})\leq\Delta(G^{2})+1\leq\Delta^{2}(G)+1$ holds for every graph $G$. So, $\chi(G^{2})=\Delta^{2}(G)+1$ is a necessary condition for $\chi_{\mathrm{so}}(G)=\Delta^{2}(G)+1$. This problem has a rich history; we refer to [5] for the information needed here.
It is well known that a connected graph $G$ with $\Delta=\Delta(G)>2$ has $\chi(G^{2})=\Delta^{2}(G)+1$ if and only if $G^{2}=K_{\Delta^{2}(G)+1}$. As discovered by Hoffman and Singleton [8], such graphs exist if and only if $\Delta\in\{2,3,7\}$ and possibly $\Delta=57$ (whose existence is a major open problem in graph theory). We consider several cases.
(1) If $\Delta=1$, then clearly $G=K_{2}$ and $\chi_{\mathrm{so}}(G)=2=\Delta^{2}(G)+1$. We immediately infer that for $H=tK_{2}$, $\alpha_{\mathrm{od}}(H)=t=|H|/\chi_{\mathrm{so}}(H)=|H|/(\Delta^{2}(H)+1)$.
(2) For $\Delta=2$ we already proved in [3] that $\chi_{\mathrm{so}}(G)\leq 5$, with equality if and only if at least one component is $C_{5}$. Hence $\alpha_{\mathrm{od}}(G)\geq|G|/5=|G|/(\Delta^{2}(G)+1)$, with equality if and only if $G\cong tC_{5}$.
(3) $\Delta=3$.
It is proved in [6] that if $\Delta(G)\geq 3$ and $G$ is a connected graph other than the Petersen graph, the Hoffman–Singleton graph, or a graph with $\Delta(G)=57$ and $G^{2}=K_{57^{2}+1}$, then $\chi(G^{2})\leq\Delta^{2}(G)-1$. However, for $\chi_{\mathrm{so}}(G)=\Delta^{2}(G)+1$, the Petersen graph is excluded by Proposition 14, and the Hoffman–Singleton graph is excluded by Proposition 15. Hence it remains to consider the case $\Delta=57$, if it exists at all. To exclude this final possibility we need the following lemma.
Lemma 30.
Denote $\epsilon(\Delta)=0$ if $\Delta$ is odd and $\epsilon(\Delta)=1$ if $\Delta$ is even. If $G$ has girth at least $5$ and $\Delta(G)=\Delta$, then:
  • $(i)$
    $\alpha_{\mathrm{od}}(G)\geq\Delta-\epsilon(\Delta)$.
  • $(ii)$
    $\chi_{\mathrm{so}}(G)\leq n-(\Delta-\epsilon(\Delta))+1$.
Proof.
Assume that $G$ has girth at least $5$. Let $v$ be a vertex with $\deg(v)=\Delta(G)$. Consider $A=N(v)$, or one fewer vertex if $\Delta$ is even. Note that $A$ is independent, since $G$ has no 3-cycles. Vertex $v$ is adjacent to the entire $A$, whose size is odd. Any other vertex can have at most one neighbor in $A$, because $G$ has no 4-cycles. Hence $A$ is an odd independent set of size $\Delta-\epsilon(\Delta)$, proving $(i)$.
Coloring all vertices of $A$ with the same color and all other vertices with distinct colors, we obtain $\chi_{\mathrm{so}}\leq n-(\Delta-\epsilon(\Delta))+1$, as stated in $(ii)$. $\Box$
A Moore graph $G$ with $\Delta=57$, if it exists, is known to have girth 5. Hence, by the above lemma, $\chi_{\mathrm{so}}(G)\leq\Delta^{2}-\Delta+2$, and so it is also excluded. We conclude:
Theorem 31.
If $\Delta(G)\geq 3$, then $\alpha_{\mathrm{od}}(G)\geq|G|/((\Delta(G))^{2}-1)$.
Remark 32.
There exist some graphs with $\chi(G^{2})=\Delta^{2}(G)-1$. For $\Delta=3$ this is the critical graph for the Ramsey number $R(4,3)$ on 8 vertices, for which one can check that $\chi_{\mathrm{so}}(G)=8$ and $\alpha_{\mathrm{od}}(G)=1$. Also a further graph with these parameters exists, see Figure 1. Hence the bound in the above theorem is sharp for $\Delta=3$. There exist also such graphs with $\Delta=4$ and $\Delta=5$, mentioned in [5]; but for $\Delta\geq 6$ no example of such graphs is known so far.
Figure 1: Two 3-regular graphs of order 8
8 Concluding remarks and open problems
In this section we offer a list of open problems for future research.
We know $\alpha(G)\geq\alpha_{\mathrm{od}}(G)\geq\alpha(G^{2})$. Trivially for a connected graph $G$, $\alpha(G)=\alpha(G^{2})$ if and only if $G=K_{n}$, $n\geq 1$.
Problem 33.
  • $(i)$
    Characterize the graphs $G$ for which $\alpha_{\mathrm{od}}(G)=\alpha(G)$ holds.
  • $(ii)$
    Characterize the graphs for which $\alpha_{\mathrm{od}}(G)=\alpha(G^{2})$ holds.
The first part has been done for families like the Kneser graphs, the Cartesian products $K_{p}\,\Box\,K_{q}$, and trivially for odd regular bipartite graphs. The next assertion implies, however, that there does not exist an “induced subgraph characterization” for $\alpha_{\mathrm{od}}(G)=\alpha(G)$.
Proposition 34.
For every graph $G$ there exists a graph $H$ such that $|H|=|G|+1$, $\alpha_{\mathrm{od}}(H)=\alpha(H)=\alpha(G)+1$, and $G$ is an induced subgraph of $H$.
Proof.
Let $B$ be a largest independent set in $G$. Construct $H$ by taking a new vertex $w$ and joining it to all those vertices of $V(G)\,\diagdown\,B$ which have an even number of neighbors in $B$. (In particular, if $B$ is odd independent in $G$, then $H=G\cup K_{1}$.) By construction it follows that $B^{+}:=B\cup\{w\}$ is an odd independent set, because if $v$ has an odd number of neighbors in $B$, then its degree remains unchanged, and otherwise it is increased by 1 due to the edge $vw$. Hence $|B^{+}|\leq\alpha_{\mathrm{od}}(H)\leq\alpha(H)\leq\alpha(G)+1=|B|+1=|B^{+}|$. Thus, equality holds throughout. $\Box$
Note that the above construction of $H$ for a given $G$ needs to know a largest independent set in $G$ explicitly, which is algorithmically hard to find in general.
Problem 35.
Characterize graphs $G$ with $\chi_{\mathrm{so}}(G)=|G|$, or equivalently with $\alpha_{\mathrm{od}}(G)=1$; cf. Proposition 1.4.1 (lower bound) and 1.4.2 (upper bound).
Clearly, a necessary condition is that $G$ has diameter at most 2. One can also make the following further related observations:
  • (a)
    If $\chi_{\mathrm{so}}(G)=|G|$, then $\chi_{\mathrm{so}}(G+K_{r})=|G|+r$.
  • (b)
    If $G$ is claw-free, we know $\chi_{\mathrm{so}}(G)=\chi(G^{2})$ and $\alpha_{\mathrm{od}}(G)=\alpha(G^{2})$. Moreover, for any graph $G$ obviously $\chi(G^{2})=|G|$ holds if and only if $G$ has diameter at most $2$. Hence, for claw-free graphs $G$, $\alpha_{\mathrm{od}}(G)=1$ and $\chi_{\mathrm{so}}(G)=|G|$ if and only if $G$ has diameter at most 2.
Problem 36.
Let $r$ be odd. Characterize the connected triangle-free $r$-regular graphs for which
  • $(i)$
    $N(v)$ is a maximum odd independent set for every vertex $v$.
  • $(ii)$
    $N(v)$ is a maximum odd independent set for some vertex $v$.
  • $(iii)$
    $\alpha_{\mathrm{od}}(G)=r$.
[…]
Observe that a necessary condition for $(i)$ and $(ii)$ is to have diameter at most 3. Concerning $(iii)$ with $r\geq 3$ (not necessarily odd), since there is no request on any $N(v)$, we have $r=\alpha_{\mathrm{od}}(G)\geq|G|/(r^{2}-1)$, hence $|G|\leq r(r^{2}-1)$.
Problem 37.
Determine the computational complexity of decision problems associated with Problems 33, 35, and 36.
Problem 38.
Let $G$ be any outerplanar graph. It is known [7] that $\chi_{\mathrm{so}}(G)\leq 8$, moreover $\chi_{\mathrm{so}}(G)=7$ holds for $G=P_{6}+K_{1}$. Is it true that $\alpha_{\mathrm{od}}(G)\geq n/7$ for any outerplanar $G$ ?
The lower bound $\alpha_{\mathrm{od}}(G)\geq n/7$ is valid at least for outerplanar graphs $G$ with $\Delta(G)\leq 6$, because then also $\chi(G^{2})\leq 7$ holds, as proved in [1, Corollary 3.13].
Problem 39.
In view of the lower bounds $\alpha_{\mathrm{od}}(G)\geq|G|/(\Delta^{2}(G)+1)$ and $\alpha_{\mathrm{od}}(G)\geq|G|/(\Delta^{2}(G)-1)$ for $\Delta\geq 3$, it is of interest to consider $c(\Delta):=\min\{c:$ there exists a graph such that $\alpha_{\mathrm{od}}(G)=c|G|/(\Delta^{2}(G)+1)\}$.
Here $c(1)=c(2)=1$, but $c(3)=5/4$ realized by two graphs of order 8; one of them is the critical graph for $R(3,4)$. The graphs $K_{p}\,\Box\,K_{p}$ and $K_{p}\,\Box\,K_{p+1}$ show that we cannot expect more than $c(\Delta)=(4+o(1))\cdot\Delta$.
To finish, for hypercubes the following problem remains open.
Problem 40.
Determine $\alpha_{\mathrm{od}}(Q_{d})$ when $d$ is even, or improve upon the results in Corollary 26 and Proposition 27.
In particular, is it true that if $d$ is even, then $\lim_{d\to\infty}\alpha_{\mathrm{od}}(Q_{d})/2^{d}=1/2$ ?
Acknowledgement.
This research was supported in part by the Slovenian Research Agency ARIS program P1-0383, project J1-4351, and the annual work program of Rudolfovo (R. Š.) and by the ERC Advanced Grant “ERMiD” (Zs. T.).